OpenStudy (anonymous):

Find an expression for a rational function f(x) that satisfies the conditions: a horizontal asymptote at y = 0, vertical asymptotes at x = ±3, and contains the point (0, 1).

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (freckles):

so on the bottom we know we are going to have what factors?

OpenStudy (freckles):

x=3 will make the bottom zero so x-3 is going to be on bottom you also know x=-3 will make the bottom zero so what other factor is on bottom

OpenStudy (anonymous):

I got f (x) = -9 / x^2 - 9

OpenStudy (freckles):

\[f(0)=\frac{-9}{0-9}=\frac{-9}{-9}=1 \] yep looks good

OpenStudy (freckles):

you have the ordered pair (0,1) on f you also have the horizontal asy y=0 and the vertical asymyptotes x=-3,x=3

OpenStudy (anonymous):

Write a general formula to describe the variation: x varies inversely with the cube of y. I got x = k/y^3 for thid one

OpenStudy (freckles):

for the y=0 part we just needed the degree on bottom to be bigger than the degree on top

OpenStudy (anonymous):

oh okay

OpenStudy (freckles):

y varies inversely as the value of x can be translated as y=k/x

OpenStudy (freckles):

so you are correct

OpenStudy (anonymous):

Awesome!

OpenStudy (anonymous):

Ok I have one or two more problem